OpenStudy (anonymous):

@waterineyes write an equation of a line in slope intercept form, that is parallel to y=3x+6 and passes through the point (-10, 2.5)

OpenStudy (anonymous):

Suppose if a line l is parallel to y = 3x + 6 then the slopes of these two lines must be equal..

OpenStudy (anonymous):

yes, that makes sense

OpenStudy (anonymous):

Can you find the slope of line y = 3x + 6?? If not, then try and if not then tell me.. :)

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Wow, you are faster than me, Good.. Now our line that we want to find should have 3 as its slope.. Right??

OpenStudy (anonymous):

We have find m from one condition given.. Now we are interested in finding b or c called intercept from the other condition given.. Getting??

OpenStudy (anonymous):

If you have any difficulty anywhere then do tell me..

OpenStudy (anonymous):

*found.. Don't follow my english.. :P

OpenStudy (anonymous):

lol. ok, i am following, bu just a tiny bit confused. whats an other equation we can use to find the answer?

OpenStudy (anonymous):

What, I did not get you.. See, there are two lines in our question, one is y = 3x + 6 and other is l(say).. Now we want to find the equation of l line.. Getting??

OpenStudy (anonymous):

right ok. so how do we find be now? the equation so far looks like: y=3x+?

OpenStudy (anonymous):



OpenStudy (anonymous):

Yes you are right. we have found only m.. Now we want to find other ? mark there in my diagram.. Getting??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Following??

OpenStudy (anonymous):

Now use point slope form to find the equation, then we can convert that into slope intercept form easily.. Getting?? I will tell you how to do this, don't you worry.

OpenStudy (anonymous):

Point slope from says: If a line, x + y = c passes through a point \((x_1, y_1)\) then : \[y - y_1 = m(x - x_1)\]

OpenStudy (anonymous):

*form.

OpenStudy (anonymous):

Here I wanna tell you that point is given to you in question ie (-10, 2.5) that is nothing but \(x_1\) and \(y_1\)..

OpenStudy (anonymous):

Tell me our line l is passing through a point (-10, 2.5) or not?? If not, then read the question..

OpenStudy (anonymous):

yes its supposed to pass through it

OpenStudy (anonymous):